Estimate the value of the equilibrium constant at 510 K for each of the following reactions.

a) 2CO(g)+O2(g)⇌2CO2(g)

b) 2H2S(g)⇌2H2(g)+S2(g)

There was no additional information given.
Answer #1

a)

2CO(g)+O2(g)⇌2CO2(g)

Calculate the ΔG˚rxn for this reaction,

Calculate the ΔG˚rxn for this reaction,

ΔG˚rxn = ΔG˚product -  ΔG˚reactant

=(2 (-394.4 kJ/mole)) -(2 (-137.2 kJ/mole) + 0) = -514.4 kJ/mole = -514400 J/mole

Calculate the equilibrium constant, K

K = e-(ΔG˚rxn /RT)

= e-(-514400 /8.314 x 510)

K = 4.87 x 1052

b)

2H2S(g)⇌2H2(g)+S2(g)

Calculate the ΔG˚rxn for this reaction,

ΔG˚rxn = ΔG˚product -  ΔG˚reactant

=(2 (0)+(79.3 kJ/mole)) -(2 (-33.6 kJ/mole)) = 146.5 kJ/mole = 146500 J/mole

Calculate the equilibrium constant, K

K = e-(ΔG˚rxn /RT)

= e-(146500 J/mole /8.314 x 510)

K = 9.88 x 10-16